A North American tourist takes his 25.0 W, 120 V AC razor to Europe, finds a plug adapter, and plugs it into 240 V AC.

Assuming constant resistance, what power, in watts, does the razor consume as it is ruined?
Answer #1

Given,

P = 25 W ; V1 = 120 V

V' = 240 V

We know that, P = V^2/R

R = V^2/P = 120^2/25 = 576 Ohm

so the resistance of the razor is 576 Ohm and is constant. Now when V' = 240 V

P' = V^2/R = 240^2/576 = 100 W

Hence, P' = 100 Watts